# From Bugs to Benchmarks: A Comprehensive Survey of Software Defect Datasets

### v1

@@ 5. Availability and Usability
The availability and usability of a dataset play a crucial role for its impact on other research and practical applications. Most of the dataset papers in this study provide accompanying artifacts that range from simple download links for datasets to more complex, executable frameworks designed to automate bug reproduction or testing. There are a few instances where the actual datasets are not provided at all, or where the artifacts are not easily accessible or usable. In this section, we examine and discuss the availability and usability of the 132 datasets we survey.
@@ 5.1. Availability and Hosting Platforms
Provided Available (Timperley et al., 2024; Andrade et al., 2024; Liu et al., 2024a; Ruan et al., 2024; Ni et al., 2024; Liu et al., 2024b; Zhang et al., 2024a; Jiang et al., 2024; Cui et al., 2024; Yu et al., 2024; Xu et al., 2024; Waseem et al., 2024; Das et al., 2024; Tian et al., 2024; Jimenez et al., 2024; Pramod et al., 2024; Cai et al., 2024; Zhong et al., 2023; Avula et al., 2023; Mahbub et al., 2023; Zhao et al., 2023a; Kim and Hong, 2023; Morovati et al., 2023; Liang et al., 2023; Saavedra et al., 2024; Chen et al., 2023b; Senanayake et al., 2023; Brust et al., 2023; da Ponte et al., 2023; Guan et al., 2023; Xiong et al., 2023; Csuvik and Vidács, 2022; Pachouly et al., 2022; Liang et al., 2022; Keshavarz and Nagappan, 2022; Kim et al., 2022; Zhang et al., 2022a, b; Pereira et al., 2022; Chen et al., 2022; Paltenghi and Pradel, 2022; Wendland et al., 2021; Akimova et al., 2021; Gyimesi et al., 2021; Kim et al., 2021b; Paul et al., 2021; Zheng et al., 2021; Nikitopoulos et al., 2021; Ngo et al., 2021; Afrose et al., 2023; Kamienski et al., 2021; Makhshari and Mesbah, 2021; Ferenc et al., 2020; Sánchez et al., 2020; Wu et al., 2020; Zhang et al., 2020; Karampatsis and Sutton, 2020; Bures et al., 2020; Fan et al., 2020; Bagherzadeh et al., 2020; Wang et al., 2020; Garcia et al., 2020; Vieira et al., 2019; Benton et al., 2019; Riganelli et al., 2019; Ponta et al., 2019; Xu et al., 2019; Lin et al., 2019; Saha et al., 2018; Ferenc et al., 2018; Gkortzis et al., 2018; Zhang et al., 2018; Gao et al., 2018; Madeyski and Kawalerowicz, 2017; Tan et al., 2017; Liu et al., 2019; Lin et al., 2015; Le Goues et al., 2015, 2015; Lamkanfi et al., 2013; D’Ambros et al., 2012; Chen et al., 2021b; Selakovic and Pradel, 2017; Ho et al., 2023; Xu et al., 2023; de Santana et al., 2024; Wang et al., 2023a, b; Azad et al., 2023; Zhao et al., 2023b; Wang et al., 2023c, 2022b; Shen et al., 2021; Wang et al., 2021a; Eghbali and Pradel, 2020; Gu et al., 2019; Franco et al., 2017; Lin et al., 2017; Ye et al., 2023a; Tan et al., 2018) 100
Evolving (Silva et al., 2024; Wang et al., 2024; Lee et al., 2024; An et al., 2023; Wu et al., 2024; Jiang et al., 2023; Applis and Panichella, 2023; Bhuiyan et al., 2023; Bui et al., 2022; Song et al., 2022; Yuan et al., 2021; Widyasari et al., 2020; Tomassi et al., 2019; Radu and Nadi, 2019; Madeiral et al., 2019; Mitra and Ranganath, 2017; Just et al., 2014) 17
No Longer Available (Song et al., 2023; Du et al., 2021; Murphy et al., 2020; Ohira et al., 2015; Ku et al., 2007; Gunawi et al., 2014; Wang et al., 2017; Cifuentes et al., 2009) 8
Not Provided (Tebib et al., 2024; Yu and Wang, 2024; Liu et al., 2022; Zaman et al., 2012; Jalbert et al., 2011; Ren et al., 2020; Lu et al., 2005) 7
@@
Ideally, all datasets should be publicly available to facilitate future studies. As shown in Figure 8(a) and Table 9, we found that all but seven datasets were made publicly available at the time of paper publication. In total, 125 (125/132) dataset papers have released their datasets alongside the corresponding paper. For the 125 publicly available datasets, we further examined their availability at the time of our survey and categorized them into the following three groups:
(1) Available
@@
It means the dataset is still available at the time of our survey, as it was when the corresponding paper was published, but has not been updated. This group accounts for 100/132 of the datasets we survey.
(2) Evolving
@@
It means the dataset is publicly available, and it has been updated or evolved since the time of paper publication. Such evolution could include adding new defects, deprecating stale defects, or updating the framework or tool to facilitate the use of the dataset. This group accounts for 17/132 of the surveyed datasets.
(3) No Longer Available
@@
It means the dataset was available at some point in time, but is not available anymore as of the time of our survey. Further investigation shows the reasons include (a) institutional webpage used to host the dataset becoming no longer loadable (Ohira et al., 2015; Ku et al., 2007; Wang et al., 2017; Cifuentes et al., 2009), (b) datasets being removed from a hosting service (Song et al., 2023; Du et al., 2021), (c) unresolvable shortened links (Murphy et al., 2020), and (d) datasets being lost by accident (Gunawi et al., 2014). This group accounts for 8/132 of the surveyed datasets.
@@
Hosting platforms play a significant role in ensuring the ease of access to datasets, as our analysis of the datasets that are no longer available shows that most of them are offline due to discontinued or unstable hosting services. Figure 8(b) shows the distribution of hosting platforms for the available datasets we survey. Note that some datasets are hosted on multiple platforms. The most popular platform is GitHub, which hosts 77/117 of the available datasets. The second most popular hosting approach is to self-host a dedicated website for the dataset. Other hosting platforms, such as cloud storage providers (e.g., Google Drive, OneDrive, Box), Zenodo, and figshare, are less common but still used by some datasets.
@@ 5.2. Presentation
There are many ways to present the defects in a dataset, ranging from simply providing the link to the corresponding issue reports to creating complex executable frameworks. The way datasets present defects can impact their usability for both research and practical applications. We categorize the presentation of defects into three levels: (1) meta level, (2) code level, and (3) execution level. Table 10 shows the distribution of these presentation levels across the 117 surveyed datasets that are available or evolving.
@@
Meta Level Metadata (Andrade et al., 2024; Liu et al., 2024b; Zhang et al., 2024a; Jiang et al., 2024; Cui et al., 2024; Yu et al., 2024; Xu et al., 2024; Waseem et al., 2024; Das et al., 2024; Avula et al., 2023; Mahbub et al., 2023; Senanayake et al., 2023; da Ponte et al., 2023; Guan et al., 2023; Xiong et al., 2023; Pachouly et al., 2022; Liang et al., 2022; Keshavarz and Nagappan, 2022; Chen et al., 2022; Paltenghi and Pradel, 2022; Kim et al., 2021b; Paul et al., 2021; Nikitopoulos et al., 2021; Kamienski et al., 2021; Makhshari and Mesbah, 2021; Ferenc et al., 2020; Sánchez et al., 2020; Wu et al., 2020; Fan et al., 2020; Bagherzadeh et al., 2020; Wang et al., 2020; Garcia et al., 2020; Vieira et al., 2019; Radu and Nadi, 2019; Ponta et al., 2019; Xu et al., 2019; Madeyski and Kawalerowicz, 2017; Liu et al., 2019; Lin et al., 2015; Lamkanfi et al., 2013; D’Ambros et al., 2012; Chen et al., 2021b; Ho et al., 2023; Xu et al., 2023; de Santana et al., 2024; Wang et al., 2023a, b; Azad et al., 2023; Zhao et al., 2023b; Wang et al., 2023c, 2022b; Shen et al., 2021; Wang et al., 2021a; Eghbali and Pradel, 2020; Gu et al., 2019; Franco et al., 2017) 56
Code Level Snippet (Cai et al., 2024; Wang et al., 2024; Zhong et al., 2023; Chen et al., 2023b; Akimova et al., 2021; Zheng et al., 2021; Lin et al., 2019; Ferenc et al., 2018) 8
Diff (Kim et al., 2022; Zhang et al., 2022a; Karampatsis and Sutton, 2020; Ye et al., 2023a) 4
File (Zhao et al., 2023a; Csuvik and Vidács, 2022; Pereira et al., 2022; Ngo et al., 2021; Afrose et al., 2023; Zhang et al., 2020; Bures et al., 2020; Gao et al., 2018; Tan et al., 2017; Le Goues et al., 2015; Selakovic and Pradel, 2017) 11
Execution Level Project (Morovati et al., 2023; Zhang et al., 2018; Mitra and Ranganath, 2017; Le Goues et al., 2015) 4
Binary (Brust et al., 2023; Wendland et al., 2021; Tan et al., 2018) 3
Framework (Timperley et al., 2024; Liu et al., 2024a; Ruan et al., 2024; Ni et al., 2024; Tian et al., 2024; Jimenez et al., 2024; Pramod et al., 2024; Silva et al., 2024; Lee et al., 2024; An et al., 2023; Wu et al., 2024; Kim and Hong, 2023; Liang et al., 2023; Saavedra et al., 2024; Jiang et al., 2023; Applis and Panichella, 2023; Bhuiyan et al., 2023; Zhang et al., 2022b; Bui et al., 2022; Song et al., 2022; Gyimesi et al., 2021; Yuan et al., 2021; Widyasari et al., 2020; Benton et al., 2019; Tomassi et al., 2019; Riganelli et al., 2019; Madeiral et al., 2019; Saha et al., 2018; Gkortzis et al., 2018; Just et al., 2014; Lin et al., 2017) 31
(1) Meta Level
@@
Meta-level datasets provide the details about defects only in textual form. For example, a table listing all defects with links to corresponding issue reports is considered meta-level. Such metadata may also include the date of creation for an issue report, a severity label, or other annotations. Meta-level datasets are helpful as a reference or resource for analysis but do not include direct means for code analysis or defect reproduction. From the datasets that are still available at the time of this survey, 56/117 are meta-level.
(2) Code Level
@@
Code-level datasets provide isolated code snippets or source files from the defective program, as well as code diffs that show the changes made to the program to fix the defects. For example, a dataset that provides both buggy and fixed versions of source files without providing the entire project is considered code-level. Code-level datasets are helpful for source code analysis, but do not include direct means to run or reproduce the defects. It takes extra effort to integrate (e.g., via git apply) the provided code into a complete project to reproduce the defects. From the datasets that are still available at the time of this survey, 23/117 are code-level.
(3) Execution Level
@@
Execution-level datasets offer more hands-on resources by providing executable artifacts to help reproduce the bugs. For example, a dataset that provides a framework to automatically check out the defective program and reproduce the building and testing process is considered execution-level. Other forms of execution-level datasets include those that provide complete projects with build configurations or datasets that provide executable binaries. Execution-level datasets are particularly useful for practical applications, such as the evaluation of fault localization, test generation, or automated program repair techniques. From the datasets that are still available at the time of this survey, 38/117 are execution-level.
6. Uses of Software Defect Datasets

### v2

@@ 5. Availability and Usability (RQ3)
The availability and usability of a dataset play a crucial role for its impact on other research and practical applications. Most of the dataset papers in this study provide accompanying artifacts that range from simple download links for datasets to more complex, executable frameworks designed to automate bug reproduction or testing. There are a few instances where the actual datasets are not provided at all, or where the artifacts are not easily accessible or usable. In this section, we examine and discuss the availability and usability of the 151 datasets we survey.
@@ 5.1. Availability and Hosting Platforms
Provided Available (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/kbse/RuanLZL24; DBLP:conf/msr/NiSYZW24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:conf/msr/PramodSTSW24; DBLP:journals/jss/CaiLZZS24; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/icst/KimH23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/asiaccs/LiangYDMHZ23; DBLP:conf/icse/SaavedraSM24; DBLP:conf/raid/0001DACW23; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:journals/compsec/BrustSG23; DBLP:conf/aina/PonteRM23a; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/msr/CsuvikV22; DBLP:journals/softx/PachoulyAK22; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KeshavarzN22; DBLP:conf/msr/KimKL22; DBLP:conf/uss/ZhangP0W22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:conf/prdc/PereiraAV22; DBLP:journals/tse/ChenXLGLC22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/sigsoft/NikitopoulosDLM21; DBLP:conf/splc/NgoNNV21; DBLP:journals/tse/AfroseXRMY23; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:journals/access/SanchezDMS20; DBLP:journals/jss/WuZCWM20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/msr/KarampatsisS20; DBLP:conf/icst/BuresHA20; DBLP:conf/msr/FanL0N20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/promise/VieiraSRG19; DBLP:conf/icse/BentonGZ19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/dsa/XuYWA19; DBLP:conf/icics/LinXZX19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/promise/FerencTLSG18; DBLP:conf/msr/GkortzisMS18; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/GaoYJLYZ18; DBLP:conf/msr/MadeyskiK17; DBLP:conf/icse/TanYYMR17; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/kbse/LinMZCZ15; DBLP:journals/tse/GouesHSBDFW15; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/msr/LamkanfiPD13; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/se/SelakovicP17; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/tosem/SantanaNAA24; DBLP:journals/infsof/WangBWYGS23; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/kbse/EghbaliP20; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/FrancoGR17; DBLP:conf/oopsla/LinKCS17; DBLP:conf/kbse/YeCG23; DBLP:conf/icse/TanDGR18; DBLP:conf/fie/SandersWA24; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/icse/HuZLYH024; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/OliverDAH24; DBLP:conf/wcre/ReyesGSBM24; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19; DBLP:journals/tse/ZhaoXBCL23; DBLP:conf/kbse/XuZL23; DBLP:conf/sp/Dolan-GavittHKL16) 115
Evolving (DBLP:conf/msr/SilvaSM24; DBLP:conf/icse/WangHGWC024; DBLP:conf/icst/LeeKYY24; DBLP:conf/kbse/AnKCYY23; DBLP:conf/sigsoft/WuLZ024; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/msr/BuiSF22; DBLP:conf/issta/Song0NW0DM22; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RaduN19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/promise/MitraR17; DBLP:conf/issta/JustJE14; DBLP:conf/msr/AmannNNNM16) 18
No Longer Available (DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/issre/Du0MZ21; DBLP:conf/hotsos/MurphyBSR20; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/kbse/KuHCL07; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/kbse/WangDGGQYW17; BegBunch; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/sigsoft/Zhang0C0Z19) 10
Not Provided (DBLP:conf/pst/TebibAAG24; DBLP:conf/qrs/YuW24; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/ZamanAH12; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/iwpc/RenL020; lu2005bugbench; DBLP:conf/icse-apr/YadavW24) 8
@@
Ideally, all datasets should be publicly available to facilitate future studies. As shown in Figure 9(a) and Table 10, we found that all but eight datasets were made publicly available at the time of paper publication. In total, 143 (143/150) dataset papers have released their datasets alongside the corresponding paper. For the 143 publicly available datasets, we further examined their availability at the time of our survey and categorized them into the following three groups:
@@
(1) Available means the dataset is still available at the time of our survey, as it was when the corresponding paper was published, but has not been updated. This group accounts for 115/151 of the datasets we survey.
@@
(2) Evolving means the dataset is publicly available, and it has been updated or evolved since the time of paper publication. Such evolution could include adding new defects, deprecating stale defects, or updating the framework or tool to facilitate the use of the dataset. This group accounts for 18/151 of the surveyed datasets.
@@
(3) No Longer Available means the dataset was available at some point in time, but is not available anymore as of the time of our survey. Further investigation shows the reasons include (a) institutional webpage used to host the dataset becoming no longer loadable (DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/kbse/KuHCL07; DBLP:conf/kbse/WangDGGQYW17; BegBunch), (b) datasets being removed from a hosting service (DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/issre/Du0MZ21), (c) unresolvable shortened links (DBLP:conf/hotsos/MurphyBSR20), and (d) datasets being lost by accident (DBLP:conf/cloud/GunawiHLPDAELLM14). This group accounts for 10/151 of the surveyed datasets.
@@
Hosting platforms play a significant role in ensuring the ease of access to datasets, as our analysis of the datasets that are no longer available shows that most of them are offline due to discontinued or unstable hosting services. Figure 9(b) shows the distribution of hosting platforms for the available datasets we survey. Note that some datasets are hosted on multiple platforms. The most popular hosting platform is GitHub, which hosts 86 (86/143) available datasets. The second most popular hosting approach is to self-host a dedicated website for the dataset. Other hosting platforms, such as Zenodo, cloud storage services (e.g., Google Drive, OneDrive, Box), are also frequently used, while other platforms such as figshare, Docker Hub, GitLab, Bitbucket, and Mendeley are less common.
@@ 5.2. Presentation
There are many ways to present the defects in a dataset, ranging from simply providing the link to the corresponding issue reports to creating complex executable frameworks. The way datasets present defects can impact their usability for both research and practical applications. We categorize the presentation of defects into three levels: (1) meta level, (2) code level, and (3) execution level. Table 11 shows the distribution of these presentation levels across the 133 surveyed datasets that are available or evolving.
@@
Meta Level Metadata (DBLP:conf/edcc/AndradeLV24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:conf/aina/PonteRM23a; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:journals/softx/PachoulyAK22; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KeshavarzN22; DBLP:journals/tse/ChenXLGLC22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:conf/sigsoft/NikitopoulosDLM21; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:journals/access/SanchezDMS20; DBLP:journals/jss/WuZCWM20; DBLP:conf/msr/FanL0N20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/promise/VieiraSRG19; DBLP:conf/msr/RaduN19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/dsa/XuYWA19; DBLP:conf/msr/MadeyskiK17; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/kbse/LinMZCZ15; DBLP:conf/msr/LamkanfiPD13; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/tosem/SantanaNAA24; DBLP:journals/infsof/WangBWYGS23; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/kbse/EghbaliP20; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/FrancoGR17; DBLP:conf/fie/SandersWA24; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/icse/HuZLYH024; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/OliverDAH24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19; DBLP:journals/tse/ZhaoXBCL23) 65
Code Level Snippet (DBLP:journals/jss/CaiLZZS24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/raid/0001DACW23; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/icics/LinXZX19; DBLP:conf/promise/FerencTLSG18) 8
Diff (DBLP:conf/msr/KimKL22; DBLP:conf/uss/ZhangP0W22; DBLP:conf/msr/KarampatsisS20; DBLP:conf/kbse/YeCG23) 4
File (DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/msr/CsuvikV22; DBLP:conf/prdc/PereiraAV22; DBLP:conf/splc/NgoNNV21; DBLP:journals/tse/AfroseXRMY23; DBLP:conf/icsm/ZhangXL20; DBLP:conf/icst/BuresHA20; DBLP:conf/msr/GaoYJLYZ18; DBLP:conf/icse/TanYYMR17; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/se/SelakovicP17) 11
Execution Level Project (DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/promise/MitraR17; DBLP:journals/tse/GouesHSBDFW15) 4
Binary (DBLP:journals/compsec/BrustSG23; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/icse/TanDGR18) 3
Framework (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/kbse/RuanLZL24; DBLP:conf/msr/NiSYZW24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/msr/SilvaSM24; DBLP:conf/icst/LeeKYY24; DBLP:conf/kbse/AnKCYY23; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/icst/KimH23; DBLP:conf/asiaccs/LiangYDMHZ23; DBLP:conf/icse/SaavedraSM24; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:conf/msr/BuiSF22; DBLP:conf/issta/Song0NW0DM22; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/icse/BentonGZ19; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/msr/GkortzisMS18; DBLP:conf/issta/JustJE14; DBLP:conf/oopsla/LinKCS17; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/msr/AmannNNNM16; DBLP:conf/wcre/ReyesGSBM24; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/access/AntalVKMHF24; DBLP:conf/kbse/XuZL23; DBLP:conf/sp/Dolan-GavittHKL16) 38
@@
(1) Meta Level. Meta-level datasets provide the details about defects only in textual form. For example, a table listing all defects with links to corresponding issue reports is considered meta-level. Such metadata may also include the date of creation for an issue report, a severity label, or other annotations. Meta-level datasets are helpful as a reference or resource for analysis but do not include direct means for code analysis or defect reproduction. From the datasets that are still available at the time of this survey, 65 (65/133) are meta-level.
@@
(2) Code Level. Code-level datasets provide isolated code snippets or source files from the defective program, as well as code diffs that show the changes made to the program to fix the defects. For example, a dataset that provides both buggy and fixed versions of source files without providing the entire project is considered code-level. Code-level datasets are helpful for source code analysis, but do not include direct means to run or reproduce the defects. It takes extra effort to integrate (e.g., via git apply) the provided code into a complete project to reproduce the defects. From the datasets that are still available at the time of this survey, 23 (23/133) are code-level.
@@
(3) Execution Level. Execution-level datasets offer more hands-on resources by providing executable artifacts to help reproduce the bugs. For example, a dataset that provides a framework to automatically check out the defective program and reproduce the building and testing process is considered execution-level. Other forms of execution-level datasets include those that provide complete projects with build configurations or datasets that provide executable binaries. Execution-level datasets are particularly useful for practical applications, such as the evaluation of fault localization, test generation, or automated program repair techniques. From the datasets that are still available at the time of this survey, 45 (45/133) are execution-level.
RQ3 Summary Nearly all datasets (143/151) were released publicly at publication. Among them, 115/143 remain available, 18/143 have evolved with updates, and 10/143 are no longer available due to discontinued hosting or link decay. Long-term persistence remains a concern. GitHub dominates hosting, with 86/133 of available datasets. Dedicated websites, Zenodo, and cloud storage (e.g., Google Drive, OneDrive) are also common. Datasets that disappeared were typically hosted on institutional or self-managed servers, indicating that unstable hosting contributes to data loss. Dataset usability varies by presentation level. About 65/133 provide only meta-level information (issue links, descriptions), 23/133 include code-level artifacts (buggy files, diffs), and 45/133 offer execution-level resources (complete projects, reproduction frameworks). Execution-level datasets enable direct evaluation but require additional maintenance.
6. Uses of Software Defect Datasets (RQ4)